# A classification of $\mathbb{Q}$-linear maps from ¯×/¯$\overline{\mathbb{Q}}^{\times}/\overline{\mathbb{Q}}^{\times}_{\mathrm{tors}}$ to $\mathbb{R}$

### v1

@@ 1.1. Background
Let $\overline{\mathbb{Q}}$ be a fixed algebraic closure of $\mathbb{Q}$ and let ¯$\overline{\mathbb{Q}}^{\times}_{\mathrm{tors}}$ denote the group of roots of unity in $\overline{\mathbb{Q}}^{\times}$. We write $\mathcal{G}=\overline{\mathbb{Q}}^{\times}/\overline{\mathbb{Q}}^{\times}_{\mathrm{tors}}$ and note that $\mathcal{G}$ is a vector space over $\mathbb{Q}$ with addition and scalar multiplication given by
@@
For each number field $K$, we write $M_{K}$ to denote the set of all places of $K$. If $L/K$ is a finite extension and $w\in M_{L}$, then $w$ divides a unique place $v$ of $K$, and in this case, we write $K_{w}$ to denote the completion of $K$ with respect to $v$. Additionally, we let $p_{v}$ be the unique place of $\mathbb{Q}$ such that $v$ divides $p_{v}$ and let $\|\cdot\|_{v}$ be the unique extension to $K_{v}$ of the usual $p_{v}$-adic absolute value on $\mathbb{Q}_{v}$
@@ 1.2. Main Results
    The $\mathbb{Q}$-vector space of $\mathbb{Q}$-linear maps from $\mathcal{G}$ to $\mathbb{Q}$, i.e., the algebraic dual $\mathcal{G}^{*}$ of $\mathcal{G}$
@@
Let $\Psi(n)$ be the sum of the prime factors of $n$, counted with multiplicity (see [10, 13, 3]). Then $\Phi_{c}$ extends $\Psi$ if and only if $c(\mathbb{Q},p)=-p/(\log p)$ for all $p\neq\infty$
@@
Related to these observations, there is a version of Theorem 1.1 which classifies all continuous linear maps from $\mathcal{G}$ to , where $\mathcal{G}$ is equipped with the Weil height norm as in [2]. Specifically, if we let $\mathcal{J}^{\prime}$ be the set of consistent maps $c$ such that
$\sup\left\{\frac{c(K,v)}{\lambda(K,v)}:(K,v)\in\mathcal{J}\right\}<\infty,$
@@ 1.3. Rational Valued Linear Maps
=$\mathcal{I}^{*}=\left\{c\in\mathcal{J}^{*}:\Phi^{*}(c)\in\mathcal{G}^{*}\right\}=\left\{c\in\mathcal{J}^{*}:\Phi_{c}(\alpha)\in\mathbb{Q}\mbox{ for all }\alpha\in\mathcal{G}\right\}$
@@
We consider $K=\mathbb{Q}(\sqrt{2})$ and fix an embedding of $K$ into with ||$\cdot$|| ⋅ | denoting the usual absolute value on . We know that $K$ has two Archimedean places which we shall denote by $v_{1}$ and $v_{2}$. We may assume that
@@
for all $a,b\in\mathbb{Z}$. We observe that $K$ has class number equal to $1$, its ring of integers is $\mathcal{O}_{K}=\mathbb{Z}[\sqrt{2}]$, and $1+\sqrt{2}$ is a fundamental unit in $K$. Additionally, a prime $p\in\mathbb{Q}$ splits in $K$ if and only if $p\equiv 1,7\mod 8$.
@@
Since $\beta_{v}$ and $\beta_{w}$ are conjugates over $\mathbb{Q}$, the values in (iii) satsify $y_{v}=-y_{w}$. Therefore, we have $c(K,v)=-c(K,w)$ whenever $v$ and $w$ divide the same place of $\mathbb{Q}$, or equivalently, we have $c(\mathbb{Q},p)=0$ for all $p\in M_{\mathbb{Q}}$. The first few non-zero zero values of $c(K,v)$ are approximated in the following table.
$\begin{array}[]{|c|c|c|}p&\mbox{Factorization of }p\mbox{ in }\mathbb{Z}[\sqrt{2}]&c(K,v)\mbox{ for }v\mid p\mbox{ (approx.)}\\ \hline\cr\hline\cr\infty&\mbox{NA}&\pm 1.13459\\ \hline\cr 7&(3+\sqrt{2})(3-\sqrt{2})&\pm 0.596913\\ \hline\cr 17&(5+2\sqrt{2})(5-2\sqrt{2})&\pm 0.513516\\ \hline\cr 23&(5+\sqrt{2})(5-\sqrt{2})&\pm 0.0.513516\\ \hline\cr 31&(7+3\sqrt{2})(7-3\sqrt{2})&\pm 0.464359\\ \hline\cr 41&(7+2\sqrt{2})(7-2\sqrt{2})&\pm 0.261831\\ \hline\cr 47&(7+\sqrt{2})(7-\sqrt{2})&\pm 0.120733\\ \hline\cr 71&(11+5\sqrt{2})(11-5\sqrt{2})&\pm 0.406159\\ \hline\cr\end{array}$
@@ 2. The Kernel of $\Phi^{*}$
is called its group of units. If $K$ has $n$ Archimedean places, then Dirchlet’s Unit Theorem asserts that the $\mathcal{O}_{K}^{\times}$ has rank equal to $n-1$. If $\zeta$ is a root of unity and $\alpha_{1},\alpha_{2},\ldots,\alpha_{n-1}\in\mathcal{O}_{K}$ are such that $\{\zeta,\alpha_{1},\alpha_{2},\cdots,\alpha_{n-1}\}$ generates $\mathcal{O}_{K}^{\times}$, the then collection $\{\alpha_{1},\alpha_{2},\cdots,\alpha_{n-1}\}$ is called a fundamental set of units in $K$.
@@
For simplicity, we write $D_{i}=[K_{v_{i}},\mathbb{Q}_{v_{i}}]$ and define the following additional matrices:
@@
If $n=1$, then $K$ is an imaginary quadratic extension of $\mathbb{Q}$. Hence, we have $\lambda(K,v_{1})=1$, and the required property follows immediately from (2.2). Therefore, we shall assume that $n\geq 2$.
@@
(2.3) $AC=\begin{pmatrix}\Phi_{c}(\alpha_{1})\\ \Phi_{c}(\alpha_{2})\\ \vdots\\ \Phi_{c}(\alpha_{n})\end{pmatrix}=\begin{pmatrix}0\\ 0\\ \vdots\\ 0\end{pmatrix}$
@@
establishing that $\Phi^{*}(rc)=r\Phi^{*}(c)$ and showing that $\Phi*$ is a linear transformation.
@@ 3. Extensions of Consistent Maps
$[L:K]=\sum_{w\mid v}[L_{w}:K_{v}]\quad\mbox{and}\quad\mathrm{Norm}_{L/K}(\alpha)=\prod_{w\mid v}\mathrm{Norm}_{L_{w}/K_{v}}(\alpha)$
@@
If $L$ is any number field, let $F$ be a finite extension of $L$ containing $K$. Further, if $t$ is a place of $F$ dividing the place $v$ of $K$, then we write $y_{t}=y_{v}$. We now define the map $d_{F}:\mathcal{J}\to$ by
(3.1) $d_{F}(L,w)=\sum_{t\mid w}\frac{[F_{t}:L_{w}]}{[F:L]}y_{t}$
and claim that $d_{F}$ is independent of $F$.
To see this, we suppose that $E$ is a finite extension of $F$. In the ensuing calculations, we shall adopt the convention of writing $t$ for places of $F$ and $s$ for places of $E$. Clearly if $s\mid t$ then $y_{t}=y_{s}$, so we obtain
$d_{E}(L,w)$ $=\sum_{s\mid w}\frac{[E_{s}:L_{w}]}{[E:L]}y_{s}$
$=\sum_{t\mid w}\frac{[F_{t}:L_{w}]}{[F:L]}y_{t}\sum_{s\mid t}\frac{[E_{s}:F_{t}]}{[E:F]}$
$=\sum_{t\mid w}\frac{[F_{t}:L_{w}]}{[F:L]}\cdot y_{t}.$
Now we may define $c(L,w)=d_{F}(L,w)$, where $F$ is any number field containing both $K$ and $L$. If $L$ is a finite extension of $K$, then we apply (3.1) with $F=L$ to obtain the required property (1.5).
To prove that this consistent map is unique, we suppose that $c,d\in\mathcal{J}^{*}$ both satisfy (1.5). This means $c(L,w)=d(L,w)$ for all finite extensions $L/K$ and all places $w$ of $L$ dividing $v$. Now if $L^{\prime}$ is another number field and $w^{\prime}$ is a place of $L^{\prime}$, we may choose a number field $L$ containing both $K$ and $L^{\prime}$. Then applying the consistency of $c$ and $d$, we obtain
$c(L,w^{\prime})=\sum_{w\mid w^{\prime}}c(L,w)=\sum_{w\mid w^{\prime}}d(L,w)=d(L,w^{\prime})$
@@
Finally, we assume that $\Phi_{c}(\alpha)\in\mathbb{Q}$ for all $\alpha\in K^{\times}$. If $\beta\in\overline{\mathbb{Q}}^{\times}$ we may let $L$ be a number field containing both $\beta$ and $K$.
@@
$=\sum_{v\in M_{K}}y_{v}\sum_{w\mid v}\frac{[L_{w}:K_{v}]}{[L:K]}\log\|\beta\|_{w}$
$=\sum_{v\in M_{K}}y_{v}\sum_{w\mid v}\frac{[L_{w}:K_{v}]}{[L:K]}\log\|\mathrm{Norm}_{L_{w}/K_{v}}(\beta)\|_{v}^{1/[L_{w}:K_{v}]}$
@@
$=\frac{1}{[L:K]}\Phi(\mathrm{Norm}_{L/K}(\beta)).$
@@ 4. Surjectivity of $\Phi^{*}$
is a subgroup of $K^{\times}$ called the group of $S$-units in $K$, which according to Dirichlet’s $S$-unit Theorem (see [14, Thm. III.3.5] or [7, §1.1]), is finitely generated of rank #$\#S-1$. If $\zeta$ is a root of unity and $\alpha_{1},\alpha_{2},\ldots,\alpha_{n-1}\in U_{K,S}$ are such that $\{\zeta,\alpha_{1},\alpha_{2},\cdots,\alpha_{n-1}\}$ generates $U_{K,S}$, the then collection $\{\alpha_{1},\alpha_{2},\cdots,\alpha_{n-1}\}$ is called a fundamental set of $S$-units in $K$. Of course, this terminology generalizes our definitions from the beginning of Section 2.
@@
and note that $K_{\mathrm{div}}$ is a subgroup of $\overline{\mathbb{Q}}^{\times}$ containing ¯$\overline{\mathbb{Q}}^{\times}_{\mathrm{tors}}$. We further write
@@
and observe that $\mathcal{G}_{K}$ is a subspace of $\mathcal{G}$. If $\alpha\in K_{\mathrm{div}}$, we shall adopt the convention of writing $\bar{\alpha}$ to denote its image under the canonical homomorphism $K^{\times}\to K_{\mathrm{div}}/\overline{\mathbb{Q}}^{\times}_{\mathrm{tors}}$. For each place $v$ of $K$ and $\alpha\in K_{\mathrm{div}}$, we may assume that $\ell\in\mathbb{N}$ is such that $\alpha^{\ell}\in K$ and define
@@
    $\|\alpha^{r}\|_{v}=\|\alpha\|^{r}$ for all $\alpha\in K_{\mathrm{div}}$ and all $r\in\mathbb{Q}$.
@@
is a basis for $\mathcal{G}_{K,T_{k}}$ for all $0\leq k\leq m$. The lemma would then following by taking the special case $k=m$.
@@
    For every $\alpha\in U_{K,S}$, there exist $r_{1},r_{2},\ldots,r_{n-1}\in\mathbb{Z}$ such that $\alpha=\alpha_{1}^{r_{1}}\alpha_{2}^{r_{2}}\cdots\alpha_{n-1}^{r_{n-1}}$.
@@
Let $\bar{\alpha}\in\mathcal{G}_{K,T_{0}}$ and assume that $\alpha$ is some representative of $\bar{\alpha}$ in $K_{\mathrm{div}}$. By definition of $K_{\mathrm{div}}$, there exists $\ell\in\mathbb{N}$ such that $\alpha^{\ell}\in K^{\times}$. But clearly we also have that $\|\alpha^{\ell}\|_{w}=1$ for all non-Archimedean places $w$ of $K$, and hence, $\alpha\in U_{K,T_{0}}=U_{K,S_{\infty}}$. Then according to (i), there exist $r_{1},r_{2},\ldots,r_{n-1}\in\mathbb{Z}$ such that $\alpha^{\ell}=\alpha_{1}^{r_{1}}\alpha_{2}^{r_{2}}\cdots\alpha_{n-1}^{r_{n-1}}$. We now conclude that
@@
and proceed with the inductive step assuming that $\mathcal{B}_{k-1}$ is a basis for $\mathcal{G}_{K,T_{k-1}}$. We shall first show that $\mathcal{B}_{k}$ is linearly independent. To this end, we assume that $r_{1},r_{2},\ldots,r_{n-1},s_{1},s_{2},\ldots,s_{k}$ are such that
@@
We remind the reader that $\|\cdot\|_{w_{k}}$ is indeed a well-defined map on $\mathcal{G}_{K}$ which satisfies (A1) and (A2). Moreover, by our assumption (II), we have $\|\beta_{j}\|_{w_{k}}=1$ for all $11\leq j\leq k-$ 1. Additionally, the points $\alpha_{i}$ are ordinary units in $K$, meaning that $\|\alpha_{i}\|_{w_{k}}=1$ for all $11\leq i\leq n-$ 1. Now applying $\|\cdot\|_{w_{k}}$ to both sides of (4.1), we conclude that
@@
Again by our assumption (II), we know that $\|\bar{\beta}\|_{w_{k}}<1$, and hence, we conclude that $s_{k}=0$. Plugging this value into (4.1), we obtain
@@
so it follows from the inductive hypothesis that $r_{i}=0$ for all $11\leq i\leq n-$ 1 and $s_{j}=0$ for all $11\leq j\leq k-$ 1.
Since $\mathcal{B}_{k}$ contains $n+k-1$ elements, in order to complete the proof, it is sufficient to show that $\dim(\mathcal{G}_{K,T_{k}})\leq n+k-1$. To this end, we apply Dirichlet’s $S$-unit Theorem with $T_{k}$ in place of $S$ to obtain a fundamental set of $T_{k}$-units $\{\gamma_{1},\gamma_{2},\ldots,\gamma_{n+k-1}\}$. We claim that
@@
$\bar{\gamma}=\prod_{i=1}^{n+k-1}\gamma_{i}^{r_{i}/\ell}$
proving that $\gamma\in\mathrm{span}\{\gamma_{1},\gamma_{2},\ldots,\gamma_{n+k-1}\}$. It now follows that $\dim(\mathcal{G}_{K,T_{k}})\leq n+k-1$, establishing the lemma. ∎
@@
To prove (ii), suppose that $a,b\in(x+A)\cap(y+B)$. It follows that $a-b$ belongs to both $A$ and $B$, forcing $a-b=0$, or equivalently, $a=b$. Now (iii) follows from (i) and (ii).
@@
we see clearly that $\ker(A)=\mathrm{span}_{(}\lambda)$. We shall further write
@@
Now define the $n-1$ dimensional subspace $\Delta$ of n by
@@
Plainly we have that $\Delta\cap\ker(A)=\{0\}$ so that =${}^{n}=\Delta\oplus\ker(A)$. According to Lemma 4.3(iii), each coset of $\Delta$ shares a unique point with each coset of $\ker(A)$. Therefore, there exists a unique point ${\bf x}=(x_{1},x_{2},\ldots,x_{n})^{T}$ such that
@@
For the second statement of the Theorem, we have assumed that $\Phi_{c_{L}}(\alpha)=\Phi(\alpha)$ for all $\alpha\in L^{\times}$, so in particular, this equality holds for all $\alpha\in K^{\times}$. As a result, we have
@@
$[c_{L}-c_{K}](K,v)=[c_{L}(\mathbb{Q},\infty)-c_{K}(\mathbb{Q},\infty)]\lambda(K,v)=(r-r)\lambda(K,v)=0,$

### v2

# A classification of $\mathbb{Q}$-linear maps from $\overline{\mathbb{Q}}^{\times}/\overline{\mathbb{Q}}^{\times}_{\mathrm{tors}}$ to $\mathbb{R}$

@@ 1.1. Background
Let $\overline{\mathbb{Q}}$ be a fixed algebraic closure of $\mathbb{Q}$ and let $\overline{\mathbb{Q}}^{\times}_{\mathrm{tors}}$ denote the group of roots of unity in $\overline{\mathbb{Q}}^{\times}$. We write $\mathcal{G}=\overline{\mathbb{Q}}^{\times}/\overline{\mathbb{Q}}^{\times}_{\mathrm{tors}}$ and note that $\mathcal{G}$ is a vector space over $\mathbb{Q}$ with addition and scalar multiplication given by
@@
For each number field $K$, we write $M_{K}$ to denote the set of all places of $K$. If $L/K$ is a finite extension and $w\in M_{L}$, then $w$ divides a unique place $v$ of $K$, and in this case, we write $K_{w}$ to denote the completion of $K$ with respect to $v$. Additionally, we let $p_{v}$ be the unique place of $\mathbb{Q}$ such that $v$ divides $p_{v}$ and let $\|\cdot\|_{v}$ be the unique extension to $K_{v}$ of the usual $p_{v}$-adic absolute value on $\mathbb{Q}_{v}$. In this notation, the well-known product formula may be expressed as
$\sum_{v\in M_{K}}\frac{[K_{v}:\mathbb{Q}_{v}]}{[K:\mathbb{Q}]}\log\|\alpha\|_{v}=0$
for all non-zero elements $\alpha\in K$.
@@ 1.2. Main Results
    The $\mathbb{Q}$-vector space of $\mathbb{Q}$-linear maps from $\mathcal{G}$ to $\mathbb{Q}$, i.e., the algebraic dual $\mathcal{G}^{*}$ of $\mathcal{G}$.
@@
Let $\Psi(n)$ be the sum of the prime factors of $n$, counted with multiplicity (see [10, 13, 3]). Then $\Phi_{c}$ extends $\Psi$ if and only if $c(\mathbb{Q},p)=-p/(\log p)$ for all $p\neq\infty$.
@@
Related to these observations, there is a version of Theorem 1.1 which classifies all continuous linear maps from $\mathcal{G}$ to , where $\mathcal{G}$ is equipped with the Weil height norm as in [2]. Specifically, if we let $\mathcal{J}^{\prime}$ be the set of consistent maps $c$ for which
$\sup\left\{\left|\frac{c(K,v)}{\lambda(K,v)}\right|:(K,v)\in\mathcal{J}\right\}<\infty,$
@@ 1.3. Rational Valued Linear Maps
$\mathcal{I}^{*}=\left\{c\in\mathcal{J}^{*}:\Phi^{*}(c)\in\mathcal{G}^{*}\right\}=\left\{c\in\mathcal{J}^{*}:\Phi_{c}(\alpha)\in\mathbb{Q}\mbox{ for all }\alpha\in\mathcal{G}\right\}$
@@
We consider $K=\mathbb{Q}(\sqrt{2})$ and fix an embedding of $K$ into with $|\cdot|$ denoting the usual absolute value on . We know that $K$ has two Archimedean places which we shall denote by $v_{1}$ and $v_{2}$. We may assume that
@@
for all $a,b\in\mathbb{Z}$. We observe that $K$ has class number equal to $1$, its ring of integers is $\mathcal{O}_{K}=\mathbb{Z}[\sqrt{2}]$, and $1+\sqrt{2}$ is the fundamental unit in $K$. Additionally, a prime $p\in\mathbb{Q}$ splits in $K$ if and only if $p\equiv 1,7\mod 8$.
@@
Since $\beta_{v}$ and $\beta_{w}$ are conjugates over $\mathbb{Q}$, the values in (iii) satsify $y_{v}=-y_{w}$. Therefore, we have c(K,v)=c(K,w)c(K,v)=-c(K,w) whenever $v$ and $w$ divide the same place of $\mathbb{Q}$, or equivalently, we have $c(\mathbb{Q},p)=0$ for all $p\in M_{\mathbb{Q}}$. The first few non-zero values of $c(K,v)$ are approximated in the following table.
[…]
@@ 2. The Kernel of $\Phi^{*}$
is called its group of units. If $K$ has $n$ Archimedean places, then Dirchlet’s Unit Theorem asserts that the $\mathcal{O}_{K}^{\times}$ has rank equal to n1n-1. If $\zeta$ is a root of unity and $\alpha_{1},\alpha_{2},\ldots,\alpha_{n-1}\in\mathcal{O}_{K}$ are such that $\{\zeta,\alpha_{1},\alpha_{2},\cdots,\alpha_{n-1}\}$ generates $\mathcal{O}_{K}^{\times}$, then the collection $\{\alpha_{1},\alpha_{2},\cdots,\alpha_{n-1}\}$ is called a fundamental set of units in $K$.
@@
For simplicity, we write $D_{i}=[K_{v_{i}}:\mathbb{Q}_{v_{i}}]$ and define the following additional matrices:
@@
If $n=1$ then $\lambda(K,v_{1})=1$, and the required property follows immediately from (2.2). Therefore, we shall assume that $n\geq 2$.
@@
(2.3) $AC=\begin{pmatrix}\Phi_{c}(\alpha_{1})\\ \Phi_{c}(\alpha_{2})\\ \vdots\\ \Phi_{c}(\alpha_{n-1})\end{pmatrix}=\begin{pmatrix}0\\ 0\\ \vdots\\ 0\end{pmatrix}$
@@
establishing that $\Phi^{*}(rc)=r\Phi^{*}(c)$ and showing that $\Phi^{*}$ is a linear transformation.
@@ 3. Extensions of Consistent Maps
(3.1) $L:K]=\sum_{w\mid v}[L_{w}:K_{v}]\quad\mbox{and}\quad\mathrm{Norm}_{L/K}(\alpha)=\prod_{w\mid v}\mathrm{Norm}_{L_{w}/K_{v}}(\alpha)$
@@
If $L$ is any number field, let $F$ be a finite extension of $L$ containing $K$. Further, if $t$ is a place of $F$ dividing the place $v$ of $K$, then we write $y_{t}=y_{v}$. For each $w\in M_{L}$ we define
(3.2) $d_{F}(L,w)=\sum_{t\mid w}\frac{[F_{t}:K_{t}]}{[F:K]}y_{t},$
where the summation on the right hand side of (3.2) runs over places $t$ of $F$ dividing $w$. We claim that $d_{F}(L,w)$ is independent of $F$.
To see this, we suppose that $E$ is a finite extension of $F$. In the ensuing calculations, we shall adopt the convention of writing $t$ for places of $F$ and $s$ for places of $E$. Clearly if $s\mid t$ then $y_{t}=y_{s}$ and $K_{s}=K_{t}$, so we obtain
$d_{E}(L,w)$ $=\sum_{s\mid w}\frac{[E_{s}:K_{s}]}{[E:K]}y_{s}$
$=\sum_{t\mid w}\sum_{s\mid t}\frac{[E_{s}:F_{t}]\cdot[F_{t}:K_{t}]}{[E:F]\cdot[F:K]}y_{t}$
$=\sum_{t\mid w}\frac{[F_{t}:K_{t}]}{[F:K]}y_{t}\sum_{s\mid t}\frac{[E_{s}:F_{t}]}{[E:F]}.$
Now applying the left hand equality of (3.1), we obtain that
$d_{E}(L,w)=\sum_{t\mid w}\frac{[F_{t}:K_{t}]}{[F:K]}y_{t}=d_{F}(L,w)$
showing that $d_{F}(L,w)$ is indeed independent of $F$. Hence, we may define $c:\mathcal{J}\to$ by $c(L,w)=d_{F}(L,w)$, where $F$ is any number field containing both $K$ and $L$.
To prove that $c$ is consistent, we assume that $M$ is a finite extension of $L$ and $w$ is a place of $L$. We select a number field $F$ containing both $M$ and $K$. Then using $x$ to denote places of $M$, we obtain from (3.2) that
$\sum_{x\mid w}c(M,x)$ $=\sum_{x\mid w}d_{F}(M,x)$
$=\sum_{x\mid w}\sum_{t\mid x}\frac{[F_{t}:K_{t}]}{[F:K]}y_{t}$
$=\sum_{t\mid w}\frac{[F_{t}:K_{t}]}{[F:K]}y_{t}=d_{F}(L,w)=c(L,w),$
proving that $c$ is consistent.
To establish (1.5), we assume that $L$ is a finite extension of $K$, $v$ is a place of $K$, and $w$ is a a place of $L$ dividing $v$. Now apply (3.2) with $F=L$ to obtain
$c(L,w)=\frac{[L_{w}:K_{w}]}{[L:K]}y_{w}=\frac{[L_{w}:K_{v}]}{[L:K]}y_{v}$
which is the required property (1.5).
To prove that this consistent map is unique, we suppose that $c,d\in\mathcal{J}^{*}$ both satisfy (1.5). This means $c(L,w)=d(L,w)$ for all finite extensions $L/K$ and all places $w$ of $L$. Now if $L^{\prime}$ is an arbitrary number field and $w^{\prime}$ is a place of $L^{\prime}$, we may choose a number field $L$ containing both $K$ and $L^{\prime}$. Then applying the consistency of $c$ and $d$, we obtain
$c(L^{\prime},w^{\prime})=\sum_{w\mid w^{\prime}}c(L,w)=\sum_{w\mid w^{\prime}}d(L,w)=d(L^{\prime},w^{\prime})$
@@
Finally, we assume that $\Phi_{c}(\alpha)\in\mathbb{Q}$ for all $\alpha\in K^{\times}$. If $\beta\in\overline{\mathbb{Q}}^{\times}$ we may let $L$ be a number field containing both $\beta$ and $K$. Now applying (1.5) we obtain
@@
$=\sum_{v\in M_{K}}\sum_{w\mid v}c(L,w)\log\|\beta\|_{w}=\sum_{v\in M_{K}}y_{v}\sum_{w\mid v}\frac{[L_{w}:K_{v}]}{[L:K]}\log\|\beta\|_{w}.$
Then using the right hand equality of (3.1), we find that
$\Phi_{c}(\beta)$ $=\sum_{v\in M_{K}}y_{v}\sum_{w\mid v}\frac{[L_{w}:K_{v}]}{[L:K]}\log\|\mathrm{Norm}_{L_{w}/K_{v}}(\beta)\|_{v}^{1/[L_{w}:K_{v}]}$
@@
$=\frac{1}{[L:K]}\Phi_{c}(\mathrm{Norm}_{L/K}(\beta)).$
@@ 4. Surjectivity of $\Phi^{*}$
In order to complete the proof of Theorem 1.1, we must establish the following result.
@@
is a subgroup of $K^{\times}$ called the group of $S$-units in $K$, which according to Dirichlet’s $S$-unit Theorem (see [14, Thm. III.3.5] or [7, §1.1]), is finitely generated of rank #S1\#S-1. If $\zeta$ is a root of unity and $\alpha_{1},\alpha_{2},\ldots,\alpha_{\#S-1}\in U_{K,S}$ are such that $\{\zeta,\alpha_{1},\alpha_{2},\cdots,\alpha_{\#S-1}\}$ generates $U_{K,S}$, the then collection $\{\alpha_{1},\alpha_{2},\cdots,\alpha_{\#S-1}\}$ is called a fundamental set of $S$-units in $K$. Of course, this terminology generalizes our definitions from the beginning of Section 2.
@@
and note that $K_{\mathrm{div}}$ is a subgroup of $\overline{\mathbb{Q}}^{\times}$ containing $\overline{\mathbb{Q}}^{\times}_{\mathrm{tors}}$. We further write
@@
and observe that $\mathcal{G}_{K}$ is a subspace of $\mathcal{G}$. If $\alpha\in K_{\mathrm{div}}$, we shall adopt the convention of writing $\bar{\alpha}$ to denote its image under the canonical homomorphism $K_{\mathrm{div}}\to K_{\mathrm{div}}/\overline{\mathbb{Q}}^{\times}_{\mathrm{tors}}$. For each place $v$ of $K$ and $\alpha\in K_{\mathrm{div}}$, we may assume that $\ell\in\mathbb{N}$ is such that $\alpha^{\ell}\in K$ and define
@@
    $\|\alpha^{r}\|_{v}=\|\alpha\|_{v}^{r}$ for all $\alpha\in K_{\mathrm{div}}$ and all $r\in\mathbb{Q}$.
@@
is a basis for $\mathcal{G}_{K,T_{k}}$ for all $0\leq k\leq m$. The lemma would then follow by taking the special case $k=m$.
@@
    For every $\alpha\in U_{K,T_{0}}$, there exist $r_{1},r_{2},\ldots,r_{n-1}\in\mathbb{Z}$ such that $\alpha=\alpha_{1}^{r_{1}}\alpha_{2}^{r_{2}}\cdots\alpha_{n-1}^{r_{n-1}}$.
@@
Let $\bar{\alpha}\in\mathcal{G}_{K,T_{0}}$ and assume that $\alpha$ is some representative of $\bar{\alpha}$ in $K_{\mathrm{div}}$. By definition of $K_{\mathrm{div}}$, there exists $\ell\in\mathbb{N}$ such that $\alpha^{\ell}\in K^{\times}$. But clearly we also have that $\|\alpha^{\ell}\|_{w}=1$ for all non-Archimedean places $w$ of $K$, and hence, $\alpha^{\ell}\in U_{K,T_{0}}=U_{K,S_{\infty}}$. Then according to (i), there exist $r_{1},r_{2},\ldots,r_{n-1}\in\mathbb{Z}$ such that $\alpha^{\ell}=\alpha_{1}^{r_{1}}\alpha_{2}^{r_{2}}\cdots\alpha_{n-1}^{r_{n-1}}$. We now conclude that
@@
and proceed with the inductive step assuming that $\mathcal{B}_{k-1}$ is a basis for $\mathcal{G}_{K,T_{k-1}}$. We shall first show that $\mathcal{B}_{k}$ is linearly independent. To this end, we assume that $r_{1},r_{2},\ldots,r_{n-1},s_{1},s_{2},\ldots,s_{k}\in\mathbb{Q}$ are such that
@@
We remind the reader that $\|\cdot\|_{w_{k}}$ is indeed a well-defined map on $\mathcal{G}_{K}$ which satisfies (A1) and (A2). Moreover, by our assumption (II), we have $\|\beta_{j}\|_{w_{k}}=1$ for all $1\leq j\leq k-1$. Additionally, the points $\alpha_{i}$ are ordinary units in $K$, meaning that $\|\alpha_{i}\|_{w_{k}}=1$ for all $1\leq i\leq n-1$. Now applying $\|\cdot\|_{w_{k}}$ to both sides of (4.1), we conclude that
@@
Again by our assumption (II), we know that $\|\bar{\beta}_{k}\|_{w_{k}}<1$, and hence, we conclude that $s_{k}=0$. Plugging this value into (4.1), we obtain
@@
so it follows from the inductive hypothesis that $r_{i}=0$ for all $1\leq i\leq n-1$ and $s_{j}=0$ for all $1\leq j\leq k-1$.
Since $\mathcal{B}_{k}$ contains n+k1n+k-1 elements, in order to complete the proof, it is sufficient to show that $\dim(\mathcal{G}_{K,T_{k}})\leq n+k-1$. To this end, we apply Dirichlet’s $S$-unit Theorem with $T_{k}$ in place of $S$ to obtain a fundamental set of $T_{k}$-units $\{\gamma_{1},\gamma_{2},\ldots,\gamma_{n+k-1}\}$. We claim that
@@
$\bar{\gamma}=\prod_{i=1}^{n+k-1}\bar{\gamma}_{i}^{r_{i}/\ell}$
proving that $\bar{\gamma}\in\mathrm{span}\{\bar{\gamma}_{1},\bar{\gamma}_{2},\ldots,\bar{\gamma}_{n+k-1}\}$. It now follows that $\dim(\mathcal{G}_{K,T_{k}})\leq n+k-1$, establishing the lemma. ∎
@@
To prove (ii), suppose that $a,b\in(x+A)\cap(y+B)$. It follows that aba-b belongs to both $A$ and $B$, forcing ab=0a-b=0, or equivalently, $a=b$. Now (iii) follows from (i) and (ii).
@@
we see clearly that $\ker(A)=\{r\Lambda:r\in\}$. We shall further write
@@
Now define the n1n-1 dimensional subspace $\Delta$ of n by
@@
Plainly we have that $\Delta\cap\ker(A)=\{0\}$ so that $^{n}=\Delta\oplus\ker(A)$. According to Lemma 4.3(iii), each coset of $\Delta$ shares a unique point with each coset of $\ker(A)$. Therefore, there exists a unique point ${\bf x}=(x_{1},x_{2},\ldots,x_{n})^{T}$ such that
@@
For the second statement of the Lemma, we have assumed that $\Phi_{c_{L}}(\alpha)=\Phi(\alpha)$ for all $\alpha\in L^{\times}$, so in particular, this equality holds for all $\alpha\in K^{\times}$. As a result, we have
@@
$c_{L}-c_{K}](K,v)=[c_{L}(\mathbb{Q},\infty)-c_{K}(\mathbb{Q},\infty)]\lambda(K,v)=(r-r)\lambda(K,v)=0,$